# Minimax discrete distribution estimation with self-consumption

### v1

@@ IV-C2 Data replacement
When $k$ is fixed, we find that since $M\leq\sqrt{2}$ trivially, $m\geq\nicefrac{{c}}{{n\alpha}}$ for some $c\geq 0$. This is especially stark in contrast to the oracle-assisted infimum, which, with probability $1$, is proportional to $\nicefrac{{1}}{{n}}$, corresponding to the case where all samples are real.
V Proofs
V-A Proofs of the minimax lower bounds
@@
The lower bound argument uses Assouad’s method with the following construction, widely known in the distribution estimation literature e.g. as described in [11, 12, 13]. Consider the set of vectors $\mathcal{V}=\{+1,-1\}^{k/2}$. With each vector $v\in\mathcal{V}$, associate the distribution
@@
Assouad’s method leads to the lower bound described in the following Lemma. The full proof is provided in Appendix A for completeness.
Lemma 5
@@
Once again using $\log x\leq x-1$ gives an expression analogous to (11); the rest of the proof follows the same steps as the proof of Theorem 1
V-B Proof of the minimax upper bound
In this section, we derive an upper bound on the minimax expected loss at the $t^{\mathrm{th}}$ stage. The sequence of estimators described here is order-optimal with respect to $n_{t}$ and $\alpha_{t}$ in some regimes; finding optimal estimator sequences in other regimes remains an interesting open problem.
@@ V-B Proof of the minimax upper bound
The proof of this Lemma is deferred to Appendix A.
@@
Combining Lemmas 7 and 8, we arrive at the following intermediate result:
Lemma 9
@@
Now, $\alpha-\alpha^{2}\leq\frac{1}{\eta n_{1}}$ ensures that the RHS of (17) is a convex combination of two distributions in $\Delta_{k}$, thus ensuring that $\hat{P}^{(1)}$ is a valid distribution almost surely. Using Lemmas 7 and 8 leads to the desired upper bound on the componentwise variances. ∎
@@
Appendix A Proofs of Auxiliary Lemmas
In this section, we present complete proofs for the Lemmas and results used in Section V.
@@ Appendix A Proofs of Auxiliary Lemmas
We first prove the lemma for the $\ell_{2}^{2}$ loss. Define the sign(.)(.)( . ) function as
@@
Let $\mathcal{V}_{+j}=\{v\in\mathcal{V}:v_{j}=+1\}$ and $\mathcal{V}_{-j}=\{v\in\mathcal{V}:v_{j}=-1\}$. Note that these are sets of equal size, and their union is $\mathcal{V}$. Denote the mixtures $\frac{1}{|\mathcal{V}_{+j}|}\sum_{v\in\mathcal{V}_{+j}}\bar{\mathbf{P}}_{t,(p_{v})}$ and $\frac{1}{|\mathcal{V}_{-j}|}\sum_{v\in\mathcal{V}_{-j}}\bar{\mathbf{P}}_{t,(p_{v})}$ as $\bar{\mathbf{Q}}_{t,(+j)}$ and $\bar{\mathbf{Q}}_{t,(-j)}$ respectively. Rewriting (20) using $E_{Q}[\mathds{1}\{A\}]=Q(A)$,

### v2

@@ IV-E Data replacement
When $k$ is fixed, we find that since $M\leq\sqrt{2}$ trivially, $m\geq\nicefrac{{c}}{{n\alpha}}$ for some $c\geq 0$. This is especially stark in contrast to the oracle-assisted infimum, which, with probability $1$, is proportional to $\nicefrac{{1}}{{n}}$, corresponding to the case where all samples are real. Thus, in the data replacement setting, the ‘best-case’ minimax expected loss is of the same order as the ‘average’ oracle-assisted loss.
Appendix A Proofs of the minimax lower bounds
@@
The lower bound argument uses Assouad’s method with the following construction, widely known in the distribution estimation literature e.g. as described in [22, 23, 24, 25]. Consider the set of vectors $\mathcal{V}=\{+1,-1\}^{k/2}$. With each vector $v\in\mathcal{V}$, associate the distribution
@@
Assouad’s method leads to the lower bound described in the following Lemma. The full proof is provided in Appendix E for completeness.
Lemma 6.
@@
Once again using $\log x\leq x-1$ gives an expression analogous to (16); the rest of the proof follows the same steps as the proof of Theorem 1. ∎
Appendix B Proofs of the minimax upper bounds
In this section, we derive upper bounds on the minimax expected loss at the $t^{\mathrm{th}}$ stage.
B-A The order optimal estimator
We first describe the performance of a sequence of estimators that are order-optimal with respect to $n_{t}$ and $\alpha_{t}$ under mild conditions; finding optimal estimator sequences in other regimes remains an interesting open problem.
@@ B-A The order optimal estimator
The proof of this Lemma is deferred to Appendix E.
@@
Combining Lemmas 8 and 9, we arrive at the following intermediate result:
Lemma 10.
@@
Now, $\alpha-\alpha^{2}\leq\frac{1}{\eta n_{1}}$ ensures that the RHS of (22) is a convex combination of two distributions in $\Delta_{k}$, thus ensuring that $\hat{P}^{(1)}$ is a valid distribution almost surely. Using Lemmas 8 and 9 leads to the desired upper bound on the componentwise variances. ∎
@@
B-B The unprocessed empirical estimator sequence
@@
The proof proceeds inductively. For $t=0$, we have $E[\ell_{2}^{2}(\hat{P}_{0},p)]=\nicefrac{{1}}{{n_{0}}}$. Assume that the estimates at stages $i,0\leq i\leq t-1$ are the empirical distributions $\hat{p}_{\mathrm{emp}}(X_{0}^{n_{0}},\dots,X_{i}^{n_{i}})$ of the samples $X_{i}^{n_{i}}$ respectively, and that the error at stage $t-1$ is bounded above by $\left(\sum_{j}p[j](1-p[j])\right)\left(\sum_{i=0}^{t-1}\frac{n_{i}}{\left(\sum_{j=0}^{i}n_{j}\right)^{2}}\right)$. It is easy to inductively see that each estimator in this sequence is unbiased. When the estimator at stage $t$ is also the empirical estimator of the samples seen up to batch $t$, from (31) and (32), we have
$E[\ell_{2}^{2}(\hat{P}_{t}^{\mathrm{emp}},p)]=$ $\sum_{j\in[1:k]}E\left[\left(\frac{\sum_{i=0}^{t-1}n_{i}}{\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}}\hat{P}^{\mathrm{emp}}_{t-1}[j]+\frac{n_{t}}{\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}}\hat{p}_{\mathrm{emp}}(X_{t}^{n_{t}})[j]-p[j]\right)^{2}\right]$
$=$ $\sum_{j\in[1:k]}\Biggl{(}\left(\frac{\sum_{i=0}^{t-1}n_{i}}{\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}}\right)^{2}\cdot\text{var}\left(\hat{P}^{\mathrm{emp}}_{t-1}[j]\right)+\left(\frac{n_{t}}{\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}}\right)^{2}\cdot\text{var}\left(\hat{p}_{\mathrm{emp}}(X_{t}^{n_{t}})\right)$
$\qquad\qquad+\frac{n_{t}\cdot\sum_{i=0}^{t-1}n_{i}}{\left(\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}\right)^{2}}\cdot\text{cov}\left(\hat{P}^{\mathrm{emp}}_{t-1}[j],\hat{p}_{\mathrm{emp}}(X_{t}^{n_{t}})\right)\Biggr{)}$
$=$ $\sum_{j\in[1:k]}\Bigg{(}\left(\frac{\sum_{i=0}^{t-1}n_{i}}{\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}}\right)^{2}\cdot\text{var}\left(\hat{P}^{\mathrm{emp}}_{t-1}[j]\right)$
$\qquad\qquad+\left(\frac{n_{t}}{\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}}\right)^{2}\cdot\left(\frac{p[j](1-p[j])}{n_{t}}+\text{var}\left(\hat{P}^{\mathrm{emp}}_{t-1}[j]\right)(1-\alpha)^{2}\left(1-\frac{1}{n_{t}}\right)\right)$
$\qquad\qquad+\frac{n_{t}\cdot\sum_{i=0}^{t-1}n_{i}}{\left(\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}\right)^{2}}\cdot\left(2(1-\alpha)\text{var}\left(\hat{P}^{\mathrm{emp}}_{t-1}[j]\right)\right)\Bigg{)}$
$=$ $\sum_{j\in[1:k]}\left(\text{var}(\hat{P}_{t-1}^{\mathrm{emp}}[j])\left(\left(1-\frac{\alpha_{t}n_{t}}{\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}}\right)^{2}-\frac{(1-\alpha_{t})^{2}n_{t}}{\left(\sum_{i=0}^{t}n_{i}\right)^{2}}\right)+\frac{p[j](1-p[j])n_{t}}{\left(\sum_{i=0}^{t}n_{i}\right)^{2}}\right)$
$\leq$ $\left(\sum_{j}p[j](1-p[j])\right)\cdot\left(\sum_{i=0}^{t-1}\frac{n_{i}}{\left(\sum_{i^{\prime}=0}^{i}n_{i^{\prime}}\right)^{2}}+\frac{n_{t}}{\left(\sum_{i=0}^{t}n_{i}\right)^{2}}\right)$
$\leq$ $\left(\sum_{j}p[j](1-p[j])\right)\cdot\left(\sum_{i=0}^{t}\frac{n_{i}}{\left(\sum_{i^{\prime}=0}^{i}n_{i^{\prime}}\right)^{2}}\right).$
Taking the supremum of $\sum_{j}p[j](1-p[j])$ over $p\in\Delta_{k}$ concludes the proof. ∎
Appendix C Proofs of the oracle-assisted bounds
In this section, we prove the oracle-assisted bounds. Recall that in the oracle-assisted problem, each sample $X_{i}[s]$ is accompanied by source identity information $W_{i}[s]\sim\mathrm{Bern}(\alpha_{i})$ provided by an oracle, such that the conditional distribution of $X_{i}[s]$ is
$\mathbf{P}\left\{X_{i}[s]=j\middle|W_{i}[s]=0,\bar{\mathbf{P}}_{i-1,(p)}\right\}=$ $p[j]$
$\mathbf{P}\left\{X_{i}[s]=j\middle|W_{i}[s]=1,\bar{\mathbf{P}}_{i-1,(p)}\right\}=$ $\hat{P}_{i-1}[j].$
Proof:
Denote the conditional distribution of each sample at stage $i$ given the auxiliary information $W_{i}^{n_{i}}$ as $\tilde{P}_{X|W}(p,\hat{P}_{i-1},\alpha)$, and for each $i\in[1:n]$, denote the joint distribution of $(X_{0}^{n_{0}},W_{1}^{n_{1}},X_{1}^{n_{1}},\dots,W_{i}^{n_{i}},X_{i}^{n_{i}})$ as $\bar{\mathbf{P}}^{\prime}_{i,(p)}$. Now for any $p_{1},p_{2}\in\Delta_{k}$, for every $i\in[0:t-1]$,
$D\left(\tilde{P}_{X|W}({p_{v},\hat{P}_{i},\alpha_{i+1}})\middle\|\tilde{P}_{X|W}({p_{v-2e_{j}},\hat{P}_{i},\alpha_{i+1}})\middle|P_{W_{i+1}^{n_{i+1}}},\bar{\mathbf{P}}^{\prime}_{i,(p_{v})}\right)=\alpha D(p_{v}\|p_{v-2e_{j}}).$ (23)
With the same construction of $\mathcal{P}_{k}$ as Section A, with the chain rule of the KL divergence along with (23), we find that
$D\Big{(}\bar{\mathbf{P}}^{\prime}_{t,(p_{v})}\Big{\|}\bar{\mathbf{P}}^{\prime}_{t,(p_{v-2e_{j}})}\Big{)}=$ $n_{0}D\left(p_{v}\|p_{v-2e_{j}}\right)+\sum_{i=0}^{t-1}n_{i+1}D\left(\tilde{P}_{X|W}({p_{v},\hat{P}_{i},\alpha_{i+1}})\middle\|\tilde{P}_{X|W}({p_{v-2e_{j}},\hat{P}_{i},\alpha_{i+1}})\middle|P_{W_{i+1}^{n_{i+1}}},\bar{\mathbf{P}}^{\prime}_{i,(p_{v})}\right)$
$=$ $D\left(p_{v}\|p_{v-2e_{j}}\right)\left(\sum_{i=0}^{t}n_{i}\alpha_{i}\right)\leq\left(\sum_{i=0}^{t}n_{i}\alpha_{i}\right)\frac{8\delta^{2}}{k(1-\delta^{2})}.$
Using Lemma 6 with the choice
$\delta^{2}=\frac{k}{64\sum_{i=0}^{t}n_{i}\alpha_{i}}<1$
leads to
$r_{t,k}^{\ell_{2}^{2},\mathrm{oracle}}\geq\frac{1}{512\sum_{i=0}^{t}n_{i}\alpha_{i}}$
whenever $\sum_{i=0}^{t}n_{i}\alpha_{i}\geq\nicefrac{{k}}{{4}}$.
For the upper bound, consider the sequence of estimators $\hat{P}_{t,\mathrm{oracle}}$ that simply outputs the empirical distribution of the real samples accumulated by stage $t$ (i.e. the samples with the corresponding auxiliary $W_{i}[s]=0$), and $p_{u}$ if there are no real samples (i.e. when $W_{i}[s]=1$ for every $i,s$). Denote the number of real samples accumulated till stage $t$ as $Z_{t}$. Conditioned on $Z_{t}=z>0$, the worst-case expected loss of $\hat{P}_{t,\mathrm{oracle}}$ is bounded above as
$E\left[\ell_{2}^{2}\left(\hat{P}_{t,\mathrm{oracle}},p\right)\middle|Z_{t}=z\right]\leq\frac{1-\frac{1}{k}}{z}.$
Using the trivial upper bound of $\sqrt{2}$ on the $\ell_{2}^{2}$ loss between any two elements of $\Delta_{k}$, we have
$E\left[\ell_{2}^{2}\left(\hat{P}_{t,\mathrm{oracle}},p\right)\right]\leq\sqrt{2}\cdot\bar{\mathbf{P}}^{\prime}_{t,({p})}\left\{Z_{t}<\frac{1}{2}\sum_{i=0}^{t}n_{i}\alpha_{i}\right\}+\frac{2\left(1-\frac{1}{k}\right)}{\sum_{i=0}^{t}n_{i}\alpha_{i}}.$ (24)
Now, $E[Z_{t}]=\sum_{i=0}^{t}n_{i}\alpha_{i}$. Since $Z_{t}$ is the sum of $\sum_{i=0}^{t}n_{i}\alpha_{i}$ independent Bernoulli random variables, using the Hoeffding inequality,
$\bar{\mathbf{P}}^{\prime}_{t,({p})}\left\{Z_{t}-\sum_{i=0}^{t}n_{i}\alpha_{i}<-\frac{1}{2}\sum_{i=0}^{t}n_{i}\alpha_{i}\right\}\leq\exp\left(-\frac{1}{2}\sum_{i=0}^{t}n_{i}\alpha_{i}\right)\leq\frac{2}{\sum_{i=0}^{t}n_{i}\alpha_{i}}.$ (25)
Combining (24) and (25) completes the proof. ∎
Appendix D Proof of the Lower Bound for the Non-Matching Regime
We start with the following Lemma which will be useful in proving the proposition:
Lemma 11.
Let $a_{i},i\in[0:\infty)$, be a non-negative sequence such that $a_{t}\leq M\sum_{i=0}^{t-1}a_{i}$ for every $t$ and a universal positive constant $M$. Then
$\sum_{i=1}^{t}\frac{a_{i}}{\sqrt{\sum_{j=0}^{i-1}a_{j}}}\leq D\sqrt{\sum_{i=0}^{t}a_{i}}$ (26)
for
$D:=1+\sqrt{1+\sup_{s\geq 1}\frac{a_{s}}{\sum_{i=0}^{s-1}a_{i}}}.$
The proof of this Lemma is given in Appendix E. We are now ready to prove Proposition 3.
Proof:
The proof proceeds in two parts:
  1. 1.
    Showing that for $\alpha_{t}=\mathcal{O}\left(\nicefrac{{1}}{{\sum_{i=0}^{t-1}n_{i}\alpha_{i}^{2}}}\right)$ and $n_{t}\alpha_{t}=\mathcal{O}\left(\sum_{i=0}^{t-1}n_{i}\alpha_{i}\right)$,
    $\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}=\mathcal{O}\left({\sqrt{\sum_{i=0}^{t}n_{i}\alpha_{i}}}\right)$
  2. 2.
    Showing that when the loss of the estimates at stages $i\in[0:t-1]$ is $\mathcal{O}\left(\nicefrac{{1}}{{{\sqrt{\sum_{j=0}^{i}n_{j}\alpha_{j}}}}}\right)$, the error term
    $h_{t}(\nicefrac{{1}}{{4k}})=\mathcal{O}\left({\sqrt{\sum_{i=0}^{t}n_{i}\alpha_{i}}}\right).$
The lower bound then follows directly from Theorem 1.
We prove the first part by contradiction. Assume that $\sqrt{\sum_{i=0}^{t}n_{i}\alpha_{i}}=o\left(\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}\right)$. Then for $\epsilon>0$, there exists $t_{\epsilon}\geq 0$ such that
$\frac{1}{\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}}\leq$ $\frac{\epsilon}{\sqrt{\sum_{i=0}^{t}n_{i}\alpha_{i}}}$ $\forall t\geq t_{\epsilon}.$
Also, since $\alpha_{t}=\mathcal{O}(\nicefrac{{1}}{{\sum_{i=0}^{t-1}n_{i}\alpha_{i}^{2}}}),$ let $\alpha_{t}\leq\nicefrac{{C}}{{\sum_{i=0}^{t-1}n_{i}\alpha_{i}^{2}}}$. Then for all $t\geq t_{\epsilon}$,
$\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}\leq$ $C\sum_{i=0}^{t}\frac{n_{i}\alpha_{i}}{\sum_{j=0}^{i-1}n_{j}\alpha_{j}^{2}}$
$\leq$ $C_{\epsilon}+\epsilon C\sum_{i=0}^{t}\frac{n_{i}\alpha_{i}}{\sqrt{\sum_{j=0}^{i-1}n_{i}\alpha_{i}}}$
$\overset{(a)}{\leq}$ $C_{\epsilon}+\epsilon CD\sqrt{\sum_{i=0}^{t}n_{i}\alpha_{i}},$
where
$C_{\epsilon}=$ $\sum_{i=0}^{t_{\epsilon}}\frac{n_{i}\alpha_{i}}{\sum_{j=0}^{i-1}n_{j}\alpha_{j}^{2}}$
and $\text{and }D=$ $\left(1+\sqrt{1+\sup_{s\geq 1}\frac{n_{s}\alpha_{s}}{\sum_{i=0}^{s-1}n_{i}\alpha_{i}}}\right).$
The inequality $(a)$ is a consequence of Lemma 11. We can thus lower bound the limit
$\lim_{t\uparrow\infty}\frac{\sqrt{\sum_{i=0}^{t}n_{i}\alpha_{i}}}{\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}}\geq\frac{1}{\epsilon CD+\lim_{t\uparrow\infty}\frac{C_{\epsilon}}{\sqrt{\sum_{i=0}^{t}n_{i}\alpha_{i}}}}=\frac{1}{\epsilon CD}>0,$
resulting in a contradiction. We thus have
$\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}=\mathcal{O}\left({\sqrt{\sum_{i=0}^{t}n_{i}\alpha_{i}}}\right)$
We bound the error term as follows:
$\sum_{i=1}^{t}n_{i}\alpha_{i}g_{i}(\nicefrac{{1}}{{4k}})\overset{(b)}{\leq}$ $\sum_{i=1}^{t}n_{i}\alpha_{i}\cdot 16k^{2}\sup_{p}E[\ell_{2}^{2}(\hat{P}_{i-1},p)]$
$\leq$ $16Mk^{2}\sum_{i=1}^{t}\frac{n_{i}\alpha_{i}}{\sqrt{\sum_{j=0}^{i-1}n_{j}\alpha_{j}}}$
$\overset{(c)}{\leq}$ $16Mk^{2}D\sqrt{\sum_{i=0}^{t}n_{i}\alpha_{i}},$
where $(b)$ is a consequence of Chebyshev’s inequality and $(c)$ follows from Lemma 11 for
$D:=\left(1+\sqrt{1+\sup_{s\geq 1}\frac{n_{s}\alpha_{s}}{\sum_{i=0}^{s-1}n_{i}\alpha_{i}}}\right).$
Appendix E Proofs of Auxiliary Lemmas
In this section, we present complete proofs for the Lemmas and results used in Sections A, B, and C.
Proof:
@@ Appendix E Proofs of Auxiliary Lemmas
We first prove the lemma for the $\ell_{2}^{2}$ loss. Define the sign()($\cdot$)( ⋅ ) function as
@@
Let $\mathcal{V}_{+j}=\{v\in\mathcal{V}:v_{j}=+1\}$ and $\mathcal{V}_{-j}=\{v\in\mathcal{V}:v_{j}=-1\}$. Note that these are sets of equal size, and their union is $\mathcal{V}$. Denote the mixtures $\frac{1}{|\mathcal{V}_{+j}|}\sum_{v\in\mathcal{V}_{+j}}\bar{\mathbf{P}}_{t,(p_{v})}$ and $\frac{1}{|\mathcal{V}_{-j}|}\sum_{v\in\mathcal{V}_{-j}}\bar{\mathbf{P}}_{t,(p_{v})}$ as $\bar{\mathbf{Q}}_{t,(+j)}$ and $\bar{\mathbf{Q}}_{t,(-j)}$ respectively.
Rewriting (29) using $E_{Q}[\mathds{1}\{A\}]=Q(A)$,
@@
Proof:
We proceed by induction. It is easy to verify (26) for $t=1$. Assume that it holds for $t=u-1$. Then
$\sum_{i=1}^{u}\frac{a_{i}}{\sqrt{\sum_{j=0}^{i-1}a_{j}}}\leq$ $D\sqrt{\sum_{i=0}^{u-1}a_{i}}+\frac{a_{u}}{\sqrt{\sum_{i=0}^{u-1}a_{i}}}=\frac{D\sum_{i=0}^{u-1}a_{i}+a_{u}}{\sqrt{\sum_{i=0}^{u-1}a_{i}}}.$
Now
$\frac{D\sum_{i=0}^{u-1}a_{i}+a_{u}}{\sqrt{\sum_{i=0}^{u-1}a_{i}}}\leq$ $D\sqrt{\sum_{i=0}^{u}a_{i}}$
$\impliedby a_{u}\leq$ $D\left(\sqrt{\sum_{i=0}^{u-1}a_{i}}\sqrt{a_{u}+\sum_{i=0}^{u-1}a_{i}}-\sum_{i=0}^{u-1}a_{i}\right)$
$\impliedby a_{u}\leq$ $D\left(\sum_{i=0}^{u-1}a_{i}\right)\left(\sqrt{1+\frac{a_{u}}{\sum_{i=0}^{u-1}a_{i}}}-1\right)$
$\impliedby\frac{a_{u}}{\sum_{i=0}^{u-1}a_{i}}\leq$ $D\left(\sqrt{1+\frac{a_{u}}{\sum_{i=0}^{u-1}a_{i}}}-1\right)$
$\impliedby D\geq$ $\sqrt{1+\frac{a_{u}}{\sum_{i=0}^{u-1}a_{i}}}+1.$
Noting that $D$ is the supremum over $u$ of the RHS completes the proof. ∎